# Feynman Graph Integrals on Kähler Manifolds

### v2

@@ Appendix B Cauchy Principal Value
We modify the argument in [18, Proposition 6.6], assuming that $|f|>1$ on $\bar{B}$.
Let $h$ be a smooth function such that $h^{j_{1}}=f$ on $\bar{B}$. Define a map $\mathbb{C}^{m+n}\to\mathbb{C}^{m+n}$ by
$\zeta_{1}=z_{1}h({\mathbf{z}},\bar{\mathbf{z}}),\quad\zeta_{l}=z_{l}\quad(l=2,\dotsc,n).$
Set
$\tilde{B}=\left\{{\mathbf{z}}\in B:|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|<1\right\}.$
We may assume this map induces an isomorphism $\lambda:\tilde{B}\to B$.
Moreover, we have
$\lambda(\tilde{B}_{\delta}^{\mathbf{j}})=B_{\delta}^{\mathbf{j}},\quad\lambda(\tilde{S}_{\delta}^{\mathbf{j}})=S_{\delta}^{\mathbf{j}}.$
The inverse map $\mu:B\to\tilde{B}$ is given by
$z_{1}=\zeta_{1}\tilde{h}(\boldsymbol{\zeta},\bar{\boldsymbol{\zeta}}),\quad z_{j}=\zeta_{j}\quad(j=2,\dotsc,n),$
where $\tilde{h}$ is a smooth function on $B$ with no zeros.
Now take a function $k\in C^{\infty}(\bar{B})$. Changing coordinates via ${\mathbf{z}}=\mu(\boldsymbol{\zeta},\bar{\boldsymbol{\zeta}})$, we obtain:
$\ \ \ \ \int_{\tilde{S}_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})k({\mathbf{z}},\bar{{\mathbf{z}}})\,d\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}$ (B.9)
$=\int_{S_{\delta}^{\mathbf{j}}}{\boldsymbol{\zeta}}^{-{\mathbf{i}}}\sum_{l=0}^{i_{1}}\binom{i_{1}}{l}\left(\ln|\tilde{h}|^{2}\right)^{l}\left(\ln|\zeta_{1}|^{2}\right)^{i_{1}-l}\ln({\boldsymbol{\zeta}}_{\hat{1}},{\mathbf{k}}_{\hat{1}})\tilde{k}({\boldsymbol{\zeta}},\bar{{\boldsymbol{\zeta}}})\,d\bar{\zeta}_{1}\wedge d{\boldsymbol{\zeta}}_{\hat{1}}\wedge d\bar{{\boldsymbol{\zeta}}}_{\hat{1}}$,
@@
$\tilde{k}({\boldsymbol{\zeta}},\bar{{\boldsymbol{\zeta}}})=\tilde{h}^{-i_{1}}k(\mu,\bar{\mu})\cdot\Big{(}\bar{\tilde{h}}+\bar{\zeta}_{1}\frac{\partial\bar{\tilde{h}}}{\partial\bar{\zeta_{1}}}\Big{)}.$
It follows from B.7 and (B.9) that (B.7) holds.
To prove (B.8), we proceed by induction on $k_{1}$. Consider the case $k_{1}=0$. Thus,
[…]
$\frac{\partial b^{*}}{\partial z_{1}}=b.$
Indeed, set
$b^{*}(z):=\frac{1}{2\pi\sqrt{-1}}\int_{\mathbb{C}}\frac{b(\xi_{1}+z_{1},\bar{\xi}_{1}+\bar{z}_{1},{\mathbf{z}}_{\hat{1}},\bar{{\mathbf{z}}}_{\hat{1}})}{\bar{\xi}_{1}}\,d\xi_{1}\wedge d\bar{\xi}_{1},$ (B.10)
one can easily verify that $b^{*}$ satisfies the desired equation.
Set
$\gamma={\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b^{*}\,d\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}$. (B.11)
Then
$d\gamma={\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}.$ (B.12)
By Stokes’ theorem,
$\int_{\tilde{B}_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=\int_{\partial\tilde{B}_{\delta}^{\mathbf{j}}}\gamma=\int_{{}^{1}\tilde{B}_{\delta}^{\mathbf{j}}}\gamma-\int_{\tilde{S}_{\delta}^{\mathbf{j}}}\gamma,$ (B.13)
where
$^{1}\tilde{B}_{\delta}^{\mathbf{j}}=\Big{\{}{\mathbf{z}}\in\mathbb{C}^{m+n}:|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|=1,\sup_{a\neq 1}|z_{a}|<1,|{\mathbf{z}}_{\hat{1}}^{{\mathbf{j}}_{\hat{1}}}|>\delta\Big{\}}.$
In fact, to show the last equality of (B.13) , we use the decomposition
$\partial\tilde{B}_{\delta}^{\mathbf{j}}={}^{1}\tilde{B}_{\delta}^{\mathbf{j}}-\tilde{S}_{\delta}^{\mathbf{j}}+\sum_{l=2}^{n}{}^{l}\tilde{B}_{\delta}^{\mathbf{j}},$
with
$^{l}\tilde{B}_{\delta}^{\mathbf{j}}=\Big{\{}{\mathbf{z}}\in\mathbb{C}^{m+n}:|z_{1}h({\mathbf{z}},\bar{{\mathbf{z}}})|<1,|z_{l}|=1,\sup_{a\neq 1,l}|z_{a}|<1,|f{\mathbf{z}}^{\mathbf{j}}|>\delta\Big{\}},l\neq 1.$
It is clear that the integrals of $\gamma$ over the $^{l}\tilde{B}_{\delta}^{\mathbf{j}},l\neq 1$ vanish.
@@
$U_{\delta}=\left\{{\mathbf{z}}\in B:|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|>1,|{\mathbf{z}}_{\hat{1}}^{{\mathbf{j}}_{\hat{1}}}|>\delta\right\},$
so that
$\partial U_{\delta}=-{}^{1}\tilde{B}_{\delta}^{\mathbf{j}}+U_{\delta}^{\prime}+\sum_{l=1}^{n}{}^{l}U_{\delta},$
where
${}^{l}U_{\delta}$ $=\Big{\{}{\mathbf{z}}\in\mathbb{C}^{m+n}:|z_{l}|=1,\sup_{a\neq l}|z_{a}|<1,|{\mathbf{z}}_{\hat{l}}^{{\mathbf{j}}_{\hat{l}}}|>\delta,|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|>1\Big{\}},$
$U_{\delta}^{\prime}$ $=\left\{{\mathbf{z}}\in B:|{\mathbf{z}}_{\hat{1}}^{{\mathbf{j}}_{\hat{1}}}|=\delta,|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|>1\right\}.$
Again, the integrals of $\gamma$ over $U_{\delta}^{\prime}$ and $^{l}U_{\delta}$ for $l\neq 1$ vanish. Note that
$^{1}U_{\delta}={}^{1}B_{\delta}^{\mathbf{j}}:=\Big{\{}{\mathbf{z}}\in\mathbb{C}^{m+n}:|z_{1}|=1,\sup_{a\neq 1}|z_{a}|<1,|{\mathbf{z}}_{\hat{1}}^{{\mathbf{j}}_{\hat{1}}}|>\delta\Big{\}},$
since $|h|>1$ on $\bar{B}$, and the integral of $b$ over $U_{\delta}$ vanishes because $\mathrm{supp}(b)\subset\tilde{B}\subset\bar{B}\setminus U_{\delta}$. Hence, by Stokes’ theorem:
$0=\int_{\partial U_{\delta}}\gamma=\int_{{}^{1}\tilde{B}_{\delta}^{\mathbf{j}}}\gamma-\int_{{}^{1}B_{\delta}^{\mathbf{j}}}\gamma.$ (B.14)
Analogous to (B.13), we have
$\int_{B_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-\mathbf{i}}\ln({\mathbf{z}},\mathbf{k})bd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=\int_{{}^{1}B_{\delta}^{\mathbf{j}}}\gamma-\int_{S_{\delta}^{\mathbf{j}}}\gamma.$ (B.15)
By (B.13), (B.14) and (B.15), we obtain:
$\left(\int_{B_{\delta}^{\mathbf{j}}}-\int_{\tilde{B}_{\delta}^{\mathbf{j}}}\right){\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=\left(\int_{\tilde{S}_{\delta}^{\mathbf{j}}}-\int_{S_{\delta}^{\mathbf{j}}}\right)\gamma.$
By (B.6) and (B.7),
$\lim_{\delta\to 0}\left(\int_{\tilde{S}_{\delta}^{\mathbf{j}}}-\int_{S_{\delta}^{\mathbf{j}}}\right)\gamma=0.$
Therefore, (B.8) holds when $k_{1}=0$.
Now we consider the case $k_{1}>0$. Let $K=\mathrm{supp}(b)\subset\tilde{B}$ be the support of $b$. Then there exists $\alpha\in(0,1)$, s.t.
$K\subset K^{\prime}:=\big{\{}{\mathbf{z}}\in B:|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|\leq\alpha,\sup\{|z_{a}|:a\neq 1\}\leq\alpha\big{\}}.$
Let
$K^{\prime\prime}:=\big{\{}{\mathbf{z}}\in B:|z_{1}h({\mathbf{z}},\bar{\mathbf{z}})|\leq\alpha\big{\}}.$
[…]
We still take $b^{*}\in C^{\infty}(\bar{B})$ to be a smooth function given by (B.10) and define the (n,n1)(n,n-1)-form
$\gamma:={\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\,\phi b^{*}\,d\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}$. (B.16)
Then, (B.12) becomes
$d\gamma=\Big{(}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b+{\mathbf{z}}^{-\tilde{{\mathbf{i}}}}k_{1}(\ln|z_{1}|^{2})^{k_{1}-1}\ln({\mathbf{z}}_{\hat{1}},{\mathbf{k}}_{\hat{1}})\phi b^{*}+{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\frac{\partial\phi}{\partial{z}_{1}}b^{*}\Big{)}d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}},$ (B.17)
where $\tilde{{\mathbf{i}}}=(i_{1}+1,i_{2},\cdots.i_{m}).$ Using (B.17) and proceeding as before, we obtain
$\ \ \ \ \lim_{\delta\to 0}\left(\int_{B_{\delta}^{{\mathbf{j}}}}-\int_{\tilde{B}_{\delta}^{{\mathbf{j}}}}\right){\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$
$=-\lim_{\delta\to 0}\left(\int_{B_{\delta}^{{\mathbf{j}}}}-\int_{\tilde{B}_{\delta}^{{\mathbf{j}}}}\right){\mathbf{z}}^{-\tilde{{\mathbf{i}}}}k_{1}(\ln|z_{1}|^{2})^{k_{1}-1}\cdot\phi b^{*}\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$
$\quad-\lim_{\delta\to 0}\left(\int_{B_{\delta}^{{\mathbf{j}}}}-\int_{\tilde{B}_{\delta}^{{\mathbf{j}}}}\right){\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\cdot\frac{\partial\phi}{\partial{z}_{1}}b^{*}\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$
$=:I_{1}+I_{2}.$
[…]
Proposition B.9.
Let ${\mathbf{i}},{\mathbf{k}},{\mathbf{l}}\in\mathbb{Z}_{\geq 0}^{m}$ and ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$. Then both limits below exist and are equal:
$\lim_{\delta\rightarrow 0}\int_{B_{\delta}^{{\mathbf{j}}+{\mathbf{l}}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})bd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=\lim_{\delta\rightarrow 0}\int_{B_{\delta}^{\mathbf{j}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})bd{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$ (B.18)
for all $b\in C^{\infty}\left(\mathbb{C}^{m+n}\right)$.
Proof.
We have verified the existence of both limits. It now suffices to show that they are equal.
We proceed by induction on $m$ and $k_{1}$. When $m=1$, one observes that for each $\delta>0$, there exists $\delta^{\prime}>0$ such that $B_{\delta}^{{\mathbf{j}}+{\mathbf{l}}}=B_{\delta^{\prime}}^{{\mathbf{j}}}$. In this case, it is clear that the two limits coincide.
Now assume $m>1$ and $k_{1}=0$.
Let $\gamma$ be the form as in (B.11). As in (B.13), by Stokes’ theorem,
$\int_{B_{\delta}^{{\mathbf{j}}+{\mathbf{l}}}}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})b\,d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}=\left(\int_{{}^{1}B^{{\mathbf{j}}+{\mathbf{l}}}_{\delta}}-\int_{S_{\delta}^{{\mathbf{j}}+{\mathbf{l}}}}\right)\gamma.$ (B.19)
Moreover,
$\int_{{}^{1}B_{\delta}^{{\mathbf{j}}+{\mathbf{l}}}}\gamma=\int_{B_{\delta,\hat{1}}^{{\mathbf{j}}+{\mathbf{l}}}}{\mathbf{z}}_{\hat{1}}^{-{\mathbf{i}}_{\hat{1}}}\ln({\mathbf{z}}_{\hat{1}},{\mathbf{k}}_{\hat{1}})\,d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}\int_{|z_{1}|=1}z_{1}^{-i_{1}}b^{*}\,dz_{1}.$ (B.20)
By the induction hypothesis, the limit of this integral as $\delta\to 0$ is equal to
$\lim_{\delta\to 0}\int_{B_{\delta,\hat{1}}^{\mathbf{j}}}{\mathbf{z}}_{\hat{1}}^{-{\mathbf{i}}_{\hat{1}}}\ln({\mathbf{z}}_{\hat{1}},{\mathbf{k}}_{\hat{1}})d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}\int_{|z_{1}|=1}z_{1}^{-i_{1}}b^{*}dz_{1}=\lim_{\delta\to 0}\int_{{}^{1}B_{\delta}^{\mathbf{j}}}\gamma.$ (B.21)
Combining (B.15), (B.19)–(B.21) with (B.6), we obtain (B.18).
When $k_{1}>0$, we instead take $\gamma$ to be the form constructed as in (B.16). Then, applying the induction hypothesis as in B.8, we again obtain (B.18).
Remark B.10.
For any $\xi=(\xi_{1},\ldots,\xi_{n})\in\mathbb{C}^{n}$, define
$B(\xi):=\Big{\{}{\mathbf{z}}\in\mathbb{C}^{m+n}:\sup_{1\leq l\leq m}|z_{l}|<1,\ \sup_{1\leq l\leq n}|z_{m+l}-\xi_{l}|<1\Big{\}}\subset\mathbb{C}^{m+n}.$ (B.22)
For each ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$, define
$S^{\mathbf{j}}_{\delta}(\xi):=\{{\mathbf{z}}\in B(\xi):|{\mathbf{z}}^{\mathbf{j}}|=\delta\},\quad B_{\delta}^{\mathbf{j}}(\xi):=\{{\mathbf{z}}\in B(\xi):|{\mathbf{z}}^{\mathbf{j}}|\in(\delta,1)\}.$
Then all lemmas and propositions in this section extend to the setting of $S^{\mathbf{j}}_{\delta}(\xi)$ and $B_{\delta}^{\mathbf{j}}(\xi)$.
Proof of B.2.
We may as well assume that $|f|>1$ on $K.$
Let $K$ be the compact support of $\alpha$. Then $K$ can be covered by finitely many open sets $\{\tilde{B}_{a}\}_{a=1}^{N}$ together with an open set $V$, where each $\tilde{B}_{a}$ is a “box” centered at some point $(0,\xi)\in\mathbb{C}^{m}\times\mathbb{C}^{n}$ associated with $f$, chosen as in B.8, and $V$ is an open set satisfying
$V\cap\left\{{\mathbf{z}}\in\mathbb{C}^{n+m}:|f{\mathbf{z}}^{{\mathbf{j}}}|<\delta_{0}\right\}=\emptyset$
for some $\delta_{0}>0$. Moreover, we may assume without loss of generality that $f\neq 0$ on the closure of $\left(\bigcup_{a=1}^{N}\tilde{B}_{a}\right)\cup V.$
Let $\{\phi_{a}\}_{a=1}^{N+1}$ be a partition of unity subordinate to the open cover $\left(\bigcup_{a=1}^{N}\tilde{B}_{a}\right)\cup V$, with $\phi_{N+1}$ supported inside $V$.
It then follows from (B.8) and B.10 that
$\lim_{\delta\to 0}\sum_{a=1}^{N}\int_{|f{\mathbf{z}}^{{\mathbf{j}}}|>\delta}\phi_{a}\alpha=\lim_{\delta\to 0}\sum_{a=1}^{N}\int_{|{\mathbf{z}}^{{\mathbf{j}}}|>\delta}\phi_{a}\alpha.$ (B.23)
Since $f\neq 0$ on $K$, there exists $\delta_{0}^{\prime}>0$ such that
$V\cap\left\{{\mathbf{z}}\in\mathbb{C}^{n+m}:|{\mathbf{z}}^{{\mathbf{j}}}|<\delta_{0}^{\prime}\right\}=\emptyset.$
Then for any $\delta<\min\{\delta_{0},\delta_{0}^{\prime}\}$, we have
$\int_{|f{\mathbf{z}}^{{\mathbf{j}}}|<\delta}\phi_{N+1}\alpha=\int_{|{\mathbf{z}}^{{\mathbf{j}}}|<\delta}\phi_{N+1}\alpha.$ (B.24)
By combining (B.23) and (B.24), we conclude that
$\lim_{\delta\to 0}\int_{|f{\mathbf{z}}^{{\mathbf{j}}}|>\delta}\alpha=\lim_{\delta\to 0}\int_{|{\mathbf{z}}^{{\mathbf{j}}}|>\delta}\alpha.$ (B.25)
Now let ${\mathbf{j}}^{\prime}=(j_{1}^{\prime},\ldots,j_{m}^{\prime})\in\mathbb{Z}_{>0}^{m}$, and define
$\tilde{{\mathbf{j}}}:=\left(\max\{j_{1},j_{1}^{\prime}\},\ldots,\max\{j_{m},j_{m}^{\prime}\}\right).$
Then, by applying a similar partition of unity argument and using B.9, we obtain
$\lim_{\delta\to 0}\int_{|{\mathbf{z}}^{{\mathbf{j}}}|>\delta}\alpha=\lim_{\delta\to 0}\int_{|{\mathbf{z}}^{\tilde{{\mathbf{j}}}}|>\delta}\alpha=\lim_{\delta\to 0}\int_{|{\mathbf{z}}^{{\mathbf{j}}^{\prime}}|>\delta}\alpha.$ (B.26)
It follows from (B.25) and (B.26) that the limit in (B.1) is independent of the choices of $f$ and $\mathbf{j}$}.

### v3

@@ Appendix B Cauchy Principal Value
We may assume that $|f|>1$ on a precompact open neighborhood $K$ containing the support of $g$.
Let $h$ be a real smooth function such that $h^{j_{1}}=f$ on $K$, and set
$\Lambda:=3\sup_{{\mathbf{z}}\in K}|\nabla h|({\mathbf{z}}).$
For $1\leq j\leq m$, consider
$U_{j}:=\{{\mathbf{z}}\in\mathbb{C}^{m+n}:|z_{j}|<\Lambda^{-1}\}.$
Consider a map $\lambda_{j}:U_{j}\cap K\to\mathbb{C}^{m+n}$ by
$\zeta_{j}=z_{j}\,h({\mathbf{z}},\bar{\mathbf{z}}),\qquad\zeta_{l}=z_{l}\quad(l\neq j).$
Then the map
$\lambda_{j}:U_{j}\cap K\;\longrightarrow\;\lambda_{j}(U_{j}\cap K)$
is a diffeomorphism. (Indeed, by our construction of $U_{j}$, the Jacobian determinant of $\lambda_{j}$ is nonzero on $U_{j}$.) Suppose its inverse $\mu_{j}:\lambda_{j}(U_{j}\cap K)\to U_{j}\cap K$ is given by
$z_{j}=\zeta_{j}\,\tilde{h}_{j}(\boldsymbol{\zeta},\bar{\boldsymbol{\zeta}}),\qquad z_{l}=\zeta_{l}\quad(l\neq j),$
where $\tilde{h}_{j}$ is a smooth nowhere vanishing function on $\lambda_{j}(U_{j}\cap K)$.
Let
$\delta_{0}:=\Lambda^{-|{\mathbf{j}}|},\qquad K_{\delta_{0}}:=K\cap\{{\mathbf{z}}\in\mathbb{C}^{m+n}:|f{\mathbf{z}}^{\mathbf{j}}|\leq\delta_{0}\}.$
Then
$K_{\delta_{0}}\subset\bigcup_{j=1}^{m}U_{j}.$ (B.11)
Let $\{\phi_{j}\}_{j=1}^{m}$ be a partition of unity subordinate to $\{U_{j}\}_{j=1}^{m}$. By (B.11), if $\delta<\delta_{0}$ then
$\int_{|f{\mathbf{z}}^{\mathbf{j}}|=\delta}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})g\;d\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}=\sum_{j=1}^{m}\int_{|f{\mathbf{z}}^{\mathbf{j}}|=\delta}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\,\phi_{j}g\;d\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}$.
It therefore suffices to show that
$\lim_{\delta\to 0}\int_{|f{\mathbf{z}}^{\mathbf{j}}|=\delta}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})(\phi_{j}g)\;d\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}=0.$
We prove only the case $j=1$.
Changing variables via ${\mathbf{z}}=\mu_{1}(\boldsymbol{\zeta},\bar{\boldsymbol{\zeta}})$ yields
$\int_{|f{\mathbf{z}}^{\mathbf{j}}|=\delta}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\,(\phi_{1}g)({\mathbf{z}},\bar{\mathbf{z}})\;d\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}$ (B.12)
$=\int_{|\boldsymbol{\zeta}^{\mathbf{j}}|=\delta}{\boldsymbol{\zeta}}^{-{\mathbf{i}}}\sum_{l=0}^{i_{1}}\binom{i_{1}}{l}(\ln|\tilde{h}|^{2})^{l}(\ln|\zeta_{1}|^{2})^{i_{1}-l}\ln({\boldsymbol{\zeta}}_{\hat{1}},{\mathbf{k}}_{\hat{1}})\,\tilde{k}_{1}(\boldsymbol{\zeta},\bar{\boldsymbol{\zeta}})\ d\bar{\zeta}_{1}\wedge d\boldsymbol{\zeta}_{\hat{1}}\wedge d\bar{\boldsymbol{\zeta}}_{\hat{1}}$
$\quad+\int_{|\boldsymbol{\zeta}^{\mathbf{j}}|=\delta}{\boldsymbol{\zeta}}^{-{\mathbf{i}}}\sum_{l=0}^{i_{1}}\binom{i_{1}}{l}(\ln|\tilde{h}|^{2})^{l}(\ln|\zeta_{1}|^{2})^{i_{1}-l}\ln({\boldsymbol{\zeta}}_{\hat{1}},{\mathbf{k}}_{\hat{1}})\,\bar{\zeta}_{1}\,\tilde{k}_{2}(\boldsymbol{\zeta},\bar{\boldsymbol{\zeta}})\ d\zeta_{1}\wedge d\boldsymbol{\zeta}_{\hat{1}}\wedge d\bar{\boldsymbol{\zeta}}_{\hat{1}}$,
@@
$\tilde{k}_{1}(\boldsymbol{\zeta},\bar{\boldsymbol{\zeta}})=\tilde{h}_{1}^{-i_{1}}(\phi_{1}g)(\mu,\bar{\mu})\left(\bar{\tilde{h}}_{1}+\bar{\zeta}_{1}\frac{\partial\bar{\tilde{h}}_{1}}{\partial\bar{\zeta}_{1}}\right),\qquad\tilde{k}_{2}(\boldsymbol{\zeta},\bar{\boldsymbol{\zeta}})=\tilde{h}_{1}^{-i_{1}}(\phi_{1}g)(\mu,\bar{\mu})\left(\frac{\partial\bar{\tilde{h}}_{1}}{\partial\zeta_{1}}\right).$
By B.8 and (B.12), we conclude that
$\lim_{\delta\to 0}\int_{|f{\mathbf{z}}^{\mathbf{j}}|=\delta}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\,\phi_{1}g\;d\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}=0.$
Proposition B.10 (B.2).
Let ${\mathbf{i}},{\mathbf{k}}\in\mathbb{Z}_{\geq 0}^{m}$ and ${\mathbf{j}}\in\mathbb{Z}_{>0}^{m}$. Let $f\in C^{\infty}(\mathbb{C}^{m+n})$ be a real nowhere–vanishing function. Then the limit
$\lim_{\delta\to 0}\int_{|f{\mathbf{z}}^{\mathbf{j}}|>\delta}{\mathbf{z}}^{-{\mathbf{i}}}\,\ln({\mathbf{z}},{\mathbf{k}})\,g\;d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$ (B.13)
exists and is independent of the choice of $f$ and $\mathbf{j}$}, for every $g\in C_{c}^{\infty}(\mathbb{C}^{m+n})$.
Proof.
If $0\leq i_{l}\leq 1$ for $1\leq l\leq m$, then
${\mathbf{z}}^{-{\mathbf{i}}}\,\ln({\mathbf{z}},{\mathbf{k}})\,g$
is Lebesgue integrable on $\mathbb{C}^{m+n}$, so the limit (B.13) exists and is independent of $f$ and $\mathbf{j}$}.
We introduce a lexicographic order on $\mathbb{Z}_{\geq 0}\times\mathbb{Z}_{\geq 0}$ by
$(p_{1},p_{2})\leq(q_{1},q_{2})\quad\Longleftrightarrow\quad\big(p_{1}<q_{1}\big)\ \text{ or }\ \big(p_{1}=q_{1}\text{ and }p_{2}\leq q_{2}\big).$
Assume inductively that the proposition holds for all pairs $(|{\mathbf{i}}|,|{\mathbf{k}}|)<(p_{1},p_{2})$. We prove it for $(|{\mathbf{i}}|,|{\mathbf{k}}|)=(p_{1},p_{2})$.
The case $0\leq i_{l}\leq 1$ is already handled above, so we may assume that $i_{1}\geq 2$.
@@
$b\;=\;\frac{1}{1-i_{1}}\,{\mathbf{z}}^{-{\mathbf{i}}}z_{1}\,d\bar{z}_{1}\wedge d{\mathbf{z}}_{\hat{1}}\wedge d\bar{{\mathbf{z}}}_{\hat{1}}$.
We compute:
$\quad\int_{|f{\mathbf{z}}^{\mathbf{j}}|>\delta}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\,g\;d{\mathbf{z}}\wedge d\bar{{\mathbf{z}}}$
$=\int_{|f{\mathbf{z}}^{\mathbf{j}}|>\delta}d\left({\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\,g\,b\right)-\int_{|f{\mathbf{z}}^{\mathbf{j}}|>\delta}(\partial_{z_{1}}\ln({\mathbf{z}},{\mathbf{k}}))\,g\;dz_{1}\wedge b-\int_{|f{\mathbf{z}}^{\mathbf{j}}|>\delta}\ln({\mathbf{z}},{\mathbf{k}})\,\partial_{z_{1}}g\;dz_{1}\wedge b$
$=\int_{|f{\mathbf{z}}^{\mathbf{j}}|=\delta}{\mathbf{z}}^{-{\mathbf{i}}}\ln({\mathbf{z}},{\mathbf{k}})\,g\,b-\int_{|f{\mathbf{z}}^{\mathbf{j}}|>\delta}(\partial_{z_{1}}\ln({\mathbf{z}},{\mathbf{k}}))\,g\;dz_{1}\wedge b\;-\;\int_{|f{\mathbf{z}}^{\mathbf{j}}|>\delta}\ln({\mathbf{z}},{\mathbf{k}})\,\partial_{z_{1}}g\;dz_{1}\wedge b$
$=:I_{1}(\delta)+I_{2}(\delta)+I_{3}(\delta).$
By B.9,
$\lim_{\delta\to 0}I_{1}(\delta)=0.$
For $I_{2}(\delta)$, the integrand has $({\mathbf{i}},{\mathbf{k}}^{\prime})$–type divisorial singularities with $|{\mathbf{k}}^{\prime}|\leq|{\mathbf{k}}|-1$. By the induction hypothesis, $\lim_{\delta\to 0}I_{2}(\delta)$ exists and is independent of $f$ and $\mathbf{j}$}.
For $I_{3}(\delta)$, the integrand has $({\mathbf{i}}^{\prime},{\mathbf{k}})$–type divisorial singularities with $|{\mathbf{i}}^{\prime}|\leq|{\mathbf{i}}|-1$. Again by the induction hypothesis, $\lim_{\delta\to 0}I_{3}(\delta)$ exists and is independent of $f$ and $\mathbf{j}$}.
Since $\mathbb{Z}_{\geq 0}\times\mathbb{Z}_{\geq 0}$ is well–ordered by the above lexicographic order, the claim follows by transfinite induction. ∎